OpenStudy (anonymous):

Find P(not a 3) x 1 2 3 4 P(x) 0.30 0.40 0.15 0.15 i dont understand this question. please help

OpenStudy (shubhamsrg):

P(not a 3) =p(1) + p(2) + p(4) = 1- p(3) Just substitute values.

OpenStudy (anonymous):

so that means ?

OpenStudy (anonymous):

sorry im not good at this!

OpenStudy (shubhamsrg):

That means that you have study harder! :)

OpenStudy (anonymous):

lol i already know that lol

OpenStudy (shubhamsrg):

Then why not following it! ;)

OpenStudy (anonymous):

ifyour not going to help dont insult me lol.

OpenStudy (anonymous):

at lease help me find the answer then insult me. lol

OpenStudy (shubhamsrg):

Well leme try to help. So can you tell me value of P(1) from the table ?

OpenStudy (anonymous):

0.30

OpenStudy (shubhamsrg):

okay, so you know all p(1), p(2),p(3),p(4) right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

peter. thank you for helping me on my last problem!

OpenStudy (shubhamsrg):

Now, p(3) means we get a 3, p(not 3) would mean we get a number which is not 3

OpenStudy (shubhamsrg):

if not 3, then we can get 1,2 or 4. So our final asnwer would be p(1) + p(2) + p(4)

OpenStudy (anonymous):

add them all up and that should be my answer?

OpenStudy (shubhamsrg):

Yes there is an alternative way also, in case you know that sum of all probabilities = 1 ?

OpenStudy (anonymous):

they do

OpenStudy (shubhamsrg):

So p(not 3) would mean we get any number else, so answer would be 1-p(3) this is the same as p(1) + p(2) + p(4)

OpenStudy (anonymous):

answer was 0.85

OpenStudy (shubhamsrg):

Thats correct